OpenStudy (liv1234):

Check my answer please? Rae deposited $6,513.05 into a savings account with an interest rate 4.6% compounded twice a year, about how long will it take for the account to be worth $9,000 A.15 years, 0 months B. 7 years, 1 month <- my answer C. 14 years, 3 month D.17 years, 3 months @hartnn

OpenStudy (liv1234):

@hartnn

OpenStudy (liv1234):

@sammixboo @zepdrix

OpenStudy (liv1234):

@GracyGirl

OpenStudy (jamaljunior):

4.6% / 2 = 2.3% = amount the accout increased ever 6 months. $9,000 / $6513.05 = how much each dollar invested is now worth 9,000 / $6513.05 = 1.023^N ln(9,000 / $6513.05) = N(ln(1.023)) N = ln(9,000 / $6513.05) / (ln(1.023)) = 14.22 six month periods 14.22 / 2 = about 7 years, 1 month B.)

OpenStudy (jamaljunior):

you're correct

563blackghost (563blackghost):

Agreed. (Heyo :D)